OpenStudy (anonymous):

Let u = <-4, -3>. Find the unit vector in the direction of u, and write your answer in component form.

OpenStudy (anonymous):

@amistre64

OpenStudy (amistre64):

unit is a term that means 1 spose we have a length of rope that is 23 feet long; how long is one unit?

OpenStudy (anonymous):

1 foot

OpenStudy (amistre64):

good, and we can verify that because 23/23 = 1 if a tree is 3 meters tall, how tall is one unit?

OpenStudy (anonymous):

100cm

OpenStudy (amistre64):

maybe ... so how long is our vector?

OpenStudy (anonymous):

1 unit

OpenStudy (amistre64):

not quite 4^2 + 3^2 is not 1^2 graphically a vector represents the hypotenuse of a right triangle of sides x and y

OpenStudy (anonymous):

so 5 units

OpenStudy (amistre64):



OpenStudy (amistre64):

if its 5 units long, then what do we scale it by to get a vector that is a unit long?

OpenStudy (amistre64):

if something is 23 feet long, we divde it by 23 if something is 3 meters tall we divide it by 3 if something has a magnitude of K, then we divide it by K

OpenStudy (anonymous):

a 1/5 scale or 1:5

OpenStudy (amistre64):

... shoulda used L to keep it conformed to the drawing eh

OpenStudy (amistre64):

correct so we know u is already in a direction; we simply need to divide it by 1/5 to get its unit vector equivalent

OpenStudy (amistre64):

\[\frac15<-4,-3>~or~<\frac{-4}{5},\frac{-3}{5}>\]

OpenStudy (anonymous):

alright, thanks! I have one more related to the same section but not entirely the same

OpenStudy (anonymous):

This always confuses me, even when I go step by step using the book: Two forces with magnitudes of 100 and 50 pounds act on an object at angles of 50° and 160° respectively. Find the direction and magnitude of the resultant force. Round to two decimal places in all intermediate steps and in your final answer.

OpenStudy (amistre64):

your adding vectors translating the end result tho gets triggy show me what your process is so far

OpenStudy (anonymous):

one sec

OpenStudy (anonymous):

I can't figure it out right now. I've been logged off of my class and can't get back on. Can you help me get started so I can do it by hand?

OpenStudy (amistre64):

we have 2 vectors whose x and y parts need to expressed in cos and sin of the angles 100(x1,y1) + 50(x2,y2) 100(cos(50), sin(50)) + 50(cos(160),sin(160)) 100cos(50) +100sin(50) 50cos(160) +50sin(160) --------------------- X Y we get the resulting vector (X,Y) determine its length sqrt(X^2+Y^2) then comes the triggy part: tangent uses Y/X soo tan(a) = Y/X a = inverse tan(Y/X)

OpenStudy (anonymous):

So we need the length first?

OpenStudy (amistre64):

not really, the length is just asked for, otherwise (X,Y) is a vector sufficient for the cause. but its asking for the properties of the vector so we have to dissect it

OpenStudy (amistre64):

have you determined the (X,Y) parts yet?

OpenStudy (anonymous):

Sorry had to talk to my dad real quick. Working on it now

OpenStudy (anonymous):

So the vector would be (100cos(50) +100sin(50), 50cos(160) +50sin(160)) right? Just have to simplify it

OpenStudy (amistre64):

like parts .... add like parts cos to cos and sin to sin

OpenStudy (anonymous):

So would sin be x and cos be y?

OpenStudy (anonymous):

or cos is x and sin is y?

OpenStudy (amistre64):

well that was a 50/50 i spose no, x is by convention a cosine measure, and y is a sin measure.

OpenStudy (anonymous):

Ok so x would be 17.29?

OpenStudy (amistre64):

 x=17.2941 is what i rnded to so i agree there